OpenStudy (kittiwitti1):

At the minute I'm just stuck on c) and d). I might be able to figure out d) but c) is kinda confusing. --- A space shuttle has a mass of 2.0 x 10^6 kg. At lift-off the engines generate an upward force of 3.0 x 10^7 N. a) What is the weight of the shuttle? (2.0 x 10^7 N) b) What is the acceleration of the shuttle when launched? (5.0 m/s/s) c) The avg. acceleration of the shuttle during its 10 minute launch is 13 m/s/s. What velocity does it attain? (7.8 km/s) d) As the engines burn, the mass of the fuel becomes less and less. Assuming the force exerted by the engines remains the same, would you expect the acceleration to increase, decrease, or remain the same? Why?

OpenStudy (kittiwitti1):

For a), I used W = Fg = mg. This becomes\[\left( 2.0 \times 10^{6} \right) 9.8\]That multiplies up to 19600000 which rounds to (with 2 sig figs): \[2.0 \times 10^{7}\] --- For b) I used \[\Sigma Fy = Fa - W = ma\]which amounts to \[\left( 3.0 x 10^{7} \right) -\left( 2.0 \times 10^{7} \right) = \left( 2.0 \times 10^{6} \right)a\]This becomes \[1.0 \times 10 ^{7} = \left( 2.0 \times 10^{6} \right)a \]which then creates the following derived equation \[a = \left(1.0 \times 10 ^{7}\right) \div \left(2.0 \times 10^{6}\right)\]This solved for the value of a makes \[a = 5.0 m/s ^{2}\]

OpenStudy (kittiwitti1):

Typo in the first equation of b), and the last one (derived). \[first: \left(3.0 \times 10^{7}\right) - \left(2.0 \times 10^{7}\right) = \left(2.0 \times 10^{6}\right)a\]\[derived: \frac{ 1.0 \times 10^{7} }{ 2.0 \times 10^{6} }\]

OpenStudy (anonymous):

For c, you assume the acceleration is constant (It's not indicated anywhere it's not). So: \[a=\frac{ dv }{ dt }\] \[adt=dv\] (Separable differential equation, easy!) \[at = v+v_{0}\] It's initial velocity is cero, right? \[at = v\] Where a = m/s^2 and t = 600 s. So, after doing lots of basic and boring stuff, v is 7.8 km/s.

OpenStudy (kittiwitti1):

EDIT: At the minute I'm just stuck on letters c) and d). I might be able to figure out d) but c) is kinda confusing. In case you haven't seen the edit. Replying to your answer, what does the d stand for? o.o Also, for d) the acceleration should decrease because it's going up, which means deceleration, or\[-g = -9.8 m/s ^{2}\] If I'm wrong correct me.

OpenStudy (anonymous):

For d) Newtons second law states that: \[F = \frac{ dp }{ dt }\] Where F and p are vectors. p is momentum, which is equal to mass*velocity-

OpenStudy (anonymous):

Oh! Sorrysorrysorry. You don't know calculus. D: I thought you did. :C Let me try again then.

OpenStudy (kittiwitti1):

I'm only taking elementary physics, we haven't learned an equation with a d in it... xD

OpenStudy (anonymous):

It's the derivative. D:

OpenStudy (anonymous):

Acceleration is defined as the change of velocity with respect to time (It's basically the same as deltaV/deltat if a is constant). When the change is inmediate (Delta is veryveryvery small) you denote it with a d, instead of a delta.

OpenStudy (kittiwitti1):

Nope, haven't learned that I'm in Pre-Calculus. Wait, what? #confused o.o

OpenStudy (kittiwitti1):

are you saying it's dv instead of \[\Delta v\] ?

OpenStudy (anonymous):

It's basically the same, if the acceleration is constant. It looks prettier though.

OpenStudy (kittiwitti1):

I mean... dv instead of \[\Delta v = v _{f} - v _{i}\]

OpenStudy (anonymous):

\[a = \frac{ deltaV }{ t }=\frac{ V-V_{0} }{ t }\] Assuming V0 is cero. \[a = \frac{ V }{ t }\] \[V\]

OpenStudy (kittiwitti1):

o.o you not using a delta sign is confusing me lolol

OpenStudy (kittiwitti1):

but I got the gist of it, I guess.

OpenStudy (anonymous):

Sorry, code got all ugly. \[a = \frac{ \Delta V }{t} = \frac{ V - V_{0} }{ t }\] Assuming V0 is cero: \[a = \frac{ V }{ t }\] Therefore: \[at = V \]

OpenStudy (kittiwitti1):

This is d) right? o.o Because c) has only acceleration...

OpenStudy (anonymous):

That's c. In c they are asking for v, when you have a and t.

OpenStudy (kittiwitti1):

wtf NEVERMIND THAT stupid question

OpenStudy (kittiwitti1):

Sorry bout that.

OpenStudy (anonymous):

Don't worry. :) Have you seen momentum?

OpenStudy (kittiwitti1):

So I would get \[v = at = 13 \times 10 = 130 ...\]WHAT IS THIS

OpenStudy (kittiwitti1):

I messed up somewhere... the answer says 7.8 km/s

OpenStudy (kittiwitti1):

Oh, is it the 10 minutes? Must be 600 s then

OpenStudy (anonymous):

\[v = at = \left(13 \frac{ meters }{ second^{2} }\right) * (600 seconds) = 7800 \frac{ meters }{ second }\]

OpenStudy (kittiwitti1):

My reaction: wut I can't believe I was so stupid. Thanks so much :) I still don't exactly get d) though...

OpenStudy (anonymous):

Just remember to always be working in the same units everything (Minuts with minutes, seconds with seconds, etc). What do you know about momentum and Newton's second law?

OpenStudy (kittiwitti1):

Momentum? D: The second law concerns inertia right? The stupid question has no answer...

OpenStudy (anonymous):

Second law basically says that: \[F = ma\] But, you know that your force is constant and that your mass decreases with time. If m decreases, what does v have to do so that F remains constant?

OpenStudy (kittiwitti1):

I do know that formula... mass decreases with time? o.o v decreases because it's proportional ?

OpenStudy (anonymous):

Let's say F is 10. If m is 5. What does a have to be? What about if m is 2?

OpenStudy (kittiwitti1):

a is proportional to m, so the values decrease/increase together

OpenStudy (anonymous):

No. If F is 10 and m is 5, a is 2. (F = ma, 10 = 5*2 ) If m is 2, a is 5. (10 = 2*5) So, when you decrease your mass, your acceleration increases.

OpenStudy (kittiwitti1):

.... Oh, right. #stupidme Sorry for being an idiot.

OpenStudy (kittiwitti1):

It's 4 am over here = n=

OpenStudy (anonymous):

Which makes sense, don't you think? If you're in a car, and your driving through the highway with lots of stuff like a sofa or something inside, you're going to move slower than without it. And, yeah, I understand, It's 5 am here. ;D

OpenStudy (kittiwitti1):

o.o that's amazing, why are you still up XD I'm a regular procrastinator so this is normal for me though

OpenStudy (kittiwitti1):

Although I did take a nap from 4 to 12 o n o

OpenStudy (anonymous):

You can also think of it this way. If F is constant then, let's just say F=constant = c. So: \[F = ma = c\] \[a = \frac{ c }{ m }\] So a is inveresly proportional to m.

OpenStudy (kittiwitti1):

So that makes it decreasing?

OpenStudy (anonymous):

I live in a very noisy neighborhood, so I like to sleep when it's noisy outside and awake when it's quiet. And, yes it does. Because a is inversely proportional to m, when m decreases, a increases. And when m increases, a decreases.

OpenStudy (anonymous):

and be awake*

OpenStudy (anonymous):

since average acceleration is given to u, just use ur acceleration equation and calculate the final velocity at the end of 10 mins!

OpenStudy (anonymous):

average acceleration = change in velocity/ time taken

OpenStudy (kittiwitti1):

what .-. I've already gotten the answer for that...?

OpenStudy (anonymous):

i thought u had problems with c)

OpenStudy (kittiwitti1):

Oh, but Lessis already answered it. :)

OpenStudy (anonymous):

oh ok..so all done in this problem?

OpenStudy (kittiwitti1):

I'm having problems with d) but I gotta see what Lessis said first.

OpenStudy (kittiwitti1):

Nah, solved.

OpenStudy (kittiwitti1):

I might have other questions just saying, so if you want to wait o.o

OpenStudy (anonymous):

lolz.. close this question first ! if u are done with this!

OpenStudy (kittiwitti1):

You have a good point lols